OpenStudy (anonymous):

two spheres of the same density have the ratio of 4 to 9 in surface area. if the small sphere weighs 10kg what does the sphere weigh

OpenStudy (anonymous):

33.75

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

surface area is directly proportional to r^2 so ratio of radius=2:3 and volume of sphere is directly proportional to r^3 so weigh=10(2/3)^3